OpenStudy (anonymous):

2log2^3 = log2^6 true of false?

OpenStudy (mathmate):

Without proper parentheses, it is not clear what the expressions are. I'll give it a try. 2log(2^3)=log((2^3)^2)=log(2^6)

OpenStudy (turingtest):

use the properties\[a\log x=\log(x^a)\]and\[\large(x^a)^b=x^{ab}\](I'm interpreting the problem in the same way as mathmate: \(2\log(2^3)=\log(2^6)\) )

OpenStudy (anonymous):

so it is true

OpenStudy (turingtest):

are you asking or telling?

OpenStudy (anonymous):

when I looked at it is true

OpenStudy (anonymous):

yes